What is the solubility of Fe(OH)3 ( iron(III) hydroxide) in...

a) ...pure water?


b) ...0.0005 M FeCl3?
Answer #1

Awwwer Solublity product (ksp) of fe conta Ksp = 2.79110~39 het s is molar solublity of fe(OH)3 * Pure water - re lots o Fes(Ksp = (0.0005+s) :(35)3 since ksp is very low s<<<< 0.0005 2.79x10-39 = (0.0005) 27 53 S = 5.91810-13 molla Solublity = 5.9 1